# A Free-fermion Formulation of Two-dimensional Ising Models

### v1

@@ Abstract
Ising model is a famous toy model in condensed matter and statistical physics. In this work we present a free-fermion formulation of the two-dimensional classical Ising models on the honeycomb, triangular and Kagomé lattices. Each Ising model is studied in the case of a zero field and the case of an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$. We employ the decorated lattice technique, star-triangle transformation and weak-graph expansion method to exactly map each Ising model in both cases into an eight-vertex model on the square lattice. The resulting vertex weights are shown to satisfy the free-fermion condition. In the zero field case, each Ising model is an even free-fermion model. In the case of the imaginary field, the Ising model on the honeycomb lattice is an even free-fermion model while those on the triangular and Kagomé lattices are odd free-fermion models. The exact solution of the Kagomé lattice Ising model in the imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$ is obtained, which has not been reported in previous works. The frustrated Ising models on the triangular and Kagomé lattices in the imaginary field still exhibit a non-zero residual entropy.
@@ I Introduction
Ising model plays a crucial role in the rearch of condensed matter and statistical physics of lattice systems, especially in understanding the phase transition via statistical mechanics. It has a long history since 1920s [1]. The one-dimensional Ising model was first solved by Ising himself [2]. Kramers and Wannier introduced the transfer matrix method to rederive the exact solution in one dimension and study the two-dimensional case [3, 4]. Onsager first used the transfer matrix method to present the exact solution of the two-dimensional Ising model on the square lattice in the case without a magnetic field [5]. This famous result has been obtained from various approaches [6, 7, 8, 9, 10, 11, 12, 13, 14]. Lee and Yang proposed the solution for the case in an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$ in their work of Lee-Yang zeros [15]. This solution has also been given by a variety of different methods [16, 17, 18, 19, 20, 21, 22]. Onsager’s solution [5] has a great influence in the field of exactly solvable lattice models [23, 24, 25, 26, 27]. Since then the mathematical exact results for Ising models have shown their beauty and power, and attracted considerable attention. Two-dimensional Ising models on the honeycomb lattice [28, 29, 30, 31], the triangular lattice [32, 33, 34, 35, 36], the Kagomé lattice [37, 38], and the checkerboard lattice [39, 40] were also exactly solved. Although the Ising model in a non-zero field is a well-known unsolved problem, the case of an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$ originated in the work of Lee and Yang [15] is special and of significant interest.
@@
In this work we focus on the exact solutions of isotropic Ising models on typical two-dimensional lattices, both in the zero field and in the imaginary field. The Hamiltonian of N spins $\left\{{{s_{i}}=\pm 1,{\rm{}}i=1,\cdots,N}\right\}$ simply reads
$H=\sum\limits_{\left\langle{ij}\right\rangle}{J{s_{i}}{s_{j}}}-{H_{{\rm{ex}}}}\sum\limits_{i=1}^{N}{{s_{i}}}~{},$ (1)
@@
where the sum $\sum\limits_{\left\langle{ij}\right\rangle}{}$ is over all nearest-neighbours and the field ${H_{{\rm{ex}}}}$ can be 0 or $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$. The approach we employ is the mapping into the free-fermion model [41, 42, 43, 44, 45], which was proposed in the vertex model problem. Vertex model is also an important and intriguing issue in statistical mechanics. Studies of vertex models such as the six-vertex [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], eight-vertex [41, 42, 58, 59, 60, 61, 62, 63, 64, 65, 66, 43] and sixteen-vertex models [67, 68, 69, 70, 71, 72] have a close relation with those of Ising models. The free-fermion models are the solvable eight-vertex models which satisfy the free-fermion condition. The exact solution of a free-fermion model can be obtained in a variety of ways, such as using the S-matrix [41], the Pfaffian [42] and the transfer matrix methods [73, 74, 75, 76, 77, 78]. Usually an eight-vertex model on the square lattice is defined as either the even subcase or the odd subcase of the general sixteen-vertex model [72], as shown in Fig. 1. In this paper we adopt the notation of vertex configurations in Ref. [69] (see Fig. 1 there), and refer to the eight-vertex model of even/odd subcase satisfying the free-fermion condition as even/odd free-fermion model.
@@
There have been lots of studies connecting the Ising model with the vertex model [21, 22, 79, 80, 81, 82]. Ref. [21] introduced a rather simple and direct correspondence between the square lattice Ising spin states and the arrow configurations of the sixteen-vertex model. The value +1/1+1/-1+ 1 / - 1 of an Ising spin is naturally mapped to the direction of the arrow settled on the corresponding site. By further performing a transformation from the sixteen vertex weights to those in the even subcase, the author showed that the Ising model is equivalent with an even free-fermion model, both for the zero field case and the imginary field case. The present paper is inspired by the success of this work. The method of transforming the square lattice Ising model to the free-fermion model is extended and applied to the Ising models on the honeycomb, triangular and Kagomé lattices.
The structure of the paper is as follows. In Sec. II we briefly introduce the methods and techniques used for transforming the Ising models into the free-fermion models, specifically the decorated lattice technique [37, 83, 84], the star-triangle transformation [37, 83, 85, 84, 86] and the weak-graph expansion [87, 88, 69, 89]. In Sec. III we present the exact results of the Ising models on the honeycomb, triangular and Kagomé lattices. Each of these models is shown to be equivalent with a free-fermion model, both in the zero field and in the imaginary field. The exact solutions are thereby obtained in the free-fermion formulation. Summary is outlined in Sec. IV.
II Methods and Techniques
It is essential to intruduce the methods and techniques for transformation of lattice systems and vertex weights, before we present the main results of the Ising models. The decorated lattice technique [37, 83, 84] and the star-triangle transformation [37, 83, 85, 84, 86] help us to obtain a lattice structure, which is convenient for translating the Boltzmann factors to the vertex weights of a sixteen-vertex model. The resulting sixteen-vertex model is then transformed into an even or odd eight-vertex model by the weak-graph expansion [87, 88, 69, 89].
@@ II.1 Decorated lattice technique
Decorated lattice technique is suitable for the case when we would like to add a spin site between a pair of nearest-neighbours, keeping the partition function invariant. As shown in Fig. 2, the interaction $J$ is replaced by $J_{1}$. In our consideration, the additional spin $\tilde{s}$ is used to represent the direction of the arrow on the edge linking $s_{1}$ and $s_{2}$. This may help us to build up an effective vertex model formulation of the Ising problem. Consider the Boltzmann factors of this pair contributed to the partition function, which should be conserved after the decorated lattice transformation. That is
$e^{-\beta J{s_{1}}{s_{2}}}=A\sum\limits_{\tilde{s}=\pm 1}{{e^{-\beta{J_{1}}\tilde{s}\left({{s_{1}}+{s_{2}}}\right)}}}~{}.$ (2)
@@
$e^{-\beta J}=A\left({{e^{-2\beta{J_{1}}}}+{e^{2\beta{J_{1}}}}}\right),~{}e^{\beta J}=2A.$
Thus the interaction $J_{1}$ and the multiplier $A$ are determined
$\cosh\left({2\beta{J_{1}}}\right)={e^{-2\beta J}}~{},$ (3)
@@
Figure 2: The decorated lattice technique.
@@ II.2 Star-triangle transformation
Fig. 3 shows the star-triangle transformation. The triangle with internal interaction $J$ is replaced by a “star” consisting of an additional spin site coupled to the three vertices with an interaction $J^{\prime}$. The partition function should be invariant after the transformation, i.e., the Boltzmann factors of this unit should be conserved
$e^{-\beta J\left({{s_{1}}{s_{2}}+{s_{2}}{s_{3}}+{s_{3}}{s_{1}}}\right)}=B\sum\limits_{\tilde{s}=\pm 1}{e^{-\beta J^{\prime}\tilde{s}\left({{s_{1}}+{s_{2}}+{s_{3}}}\right)}}~{}.$ (5)
@@
$e^{-3\beta J}=B\left(e^{-3\beta J^{\prime}}+e^{3\beta J^{\prime}}\right),~{}e^{\beta J}=B\left(e^{-\beta J^{\prime}}+e^{\beta J^{\prime}}\right).$
The interaction $J^{\prime}$ and the multiplier $B$ can be easily determined from the equations above
$\frac{{\cosh\left({3\beta J^{\prime}}\right)}}{{\cosh\left({\beta J^{\prime}}\right)}}=4{\cosh^{2}}\left({\beta J^{\prime}}\right)-3={e^{-4\beta J}}~{},$ (6)
$B={\left({{e^{-4\beta J}}+3}\right)^{-1/2}}e^{\beta J}~{}.$ (7)
One can easily verify that the star-triangle transformation directly connects the partition function of the triangular lattice model with that of the honeycomb lattice model. This enables us to simplify the derivation of the results for the triangular lattice model.
Figure 3: The star-triangle transformation.
@@ II.3 Weak-graph expansion
Weak-graph expansion is a useful tool in lattice statistical problems [87, 89]. In this paper, we use the version of weak-graph expansion method proposed by Wu, which was originally designed for the even eight-vertex model [88] and later extended to the sixteen-vertex model [69]. Applying the method to the sixteen-vertex model enables us to obtain a new set of vertex weights, with the partition function invariant. The details of the method are shown in Ref. [88], and we only list the result for all sixteen rearranged weights
${{\tilde{\omega}}_{1}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}+{\omega_{3}}+{\omega_{4}}+{\omega_{5}}+{\omega_{6}}+{\omega_{7}}+{\omega_{8}}+{\omega_{9}}+{\omega_{10}}+{\omega_{11}}+{\omega_{12}}+{\omega_{13}}+{\omega_{14}}+{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{2}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}+{\omega_{3}}+{\omega_{4}}+{\omega_{5}}+{\omega_{6}}+{\omega_{7}}+{\omega_{8}}-{\omega_{9}}-{\omega_{10}}-{\omega_{11}}-{\omega_{12}}-{\omega_{13}}-{\omega_{14}}-{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{3}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}+{\omega_{3}}+{\omega_{4}}-{\omega_{5}}-{\omega_{6}}-{\omega_{7}}-{\omega_{8}}-{\omega_{9}}-{\omega_{10}}-{\omega_{11}}-{\omega_{12}}+{\omega_{13}}+{\omega_{14}}+{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{4}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}+{\omega_{3}}+{\omega_{4}}-{\omega_{5}}-{\omega_{6}}-{\omega_{7}}-{\omega_{8}}+{\omega_{9}}+{\omega_{10}}+{\omega_{11}}+{\omega_{12}}-{\omega_{13}}-{\omega_{14}}-{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{5}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}-{\omega_{3}}-{\omega_{4}}+{\omega_{5}}+{\omega_{6}}-{\omega_{7}}-{\omega_{8}}-{\omega_{9}}-{\omega_{10}}+{\omega_{11}}+{\omega_{12}}+{\omega_{13}}+{\omega_{14}}-{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{6}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}-{\omega_{3}}-{\omega_{4}}+{\omega_{5}}+{\omega_{6}}-{\omega_{7}}-{\omega_{8}}+{\omega_{9}}+{\omega_{10}}-{\omega_{11}}-{\omega_{12}}-{\omega_{13}}-{\omega_{14}}+{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{7}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}-{\omega_{3}}-{\omega_{4}}-{\omega_{5}}-{\omega_{6}}+{\omega_{7}}+{\omega_{8}}-{\omega_{9}}-{\omega_{10}}+{\omega_{11}}+{\omega_{12}}-{\omega_{13}}-{\omega_{14}}+{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{8}}=\frac{1}{4}\left({{\omega_{1}}+{\omega_{2}}-{\omega_{3}}-{\omega_{4}}-{\omega_{5}}-{\omega_{6}}+{\omega_{7}}+{\omega_{8}}+{\omega_{9}}+{\omega_{10}}-{\omega_{11}}-{\omega_{12}}+{\omega_{13}}+{\omega_{14}}-{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{9}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}+{\omega_{3}}-{\omega_{4}}+{\omega_{5}}-{\omega_{6}}+{\omega_{7}}-{\omega_{8}}+{\omega_{9}}-{\omega_{10}}+{\omega_{11}}-{\omega_{12}}-{\omega_{13}}+{\omega_{14}}+{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{10}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}+{\omega_{3}}-{\omega_{4}}+{\omega_{5}}-{\omega_{6}}+{\omega_{7}}-{\omega_{8}}-{\omega_{9}}+{\omega_{10}}-{\omega_{11}}+{\omega_{12}}+{\omega_{13}}-{\omega_{14}}-{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{11}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}+{\omega_{3}}-{\omega_{4}}-{\omega_{5}}+{\omega_{6}}-{\omega_{7}}+{\omega_{8}}+{\omega_{9}}-{\omega_{10}}+{\omega_{11}}-{\omega_{12}}+{\omega_{13}}-{\omega_{14}}-{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{12}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}+{\omega_{3}}-{\omega_{4}}-{\omega_{5}}+{\omega_{6}}-{\omega_{7}}+{\omega_{8}}-{\omega_{9}}+{\omega_{10}}-{\omega_{11}}+{\omega_{12}}-{\omega_{13}}+{\omega_{14}}+{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{13}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}-{\omega_{3}}+{\omega_{4}}-{\omega_{5}}+{\omega_{6}}+{\omega_{7}}-{\omega_{8}}-{\omega_{9}}+{\omega_{10}}+{\omega_{11}}-{\omega_{12}}+{\omega_{13}}-{\omega_{14}}+{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{14}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}-{\omega_{3}}+{\omega_{4}}-{\omega_{5}}+{\omega_{6}}+{\omega_{7}}-{\omega_{8}}+{\omega_{9}}-{\omega_{10}}-{\omega_{11}}+{\omega_{12}}-{\omega_{13}}+{\omega_{14}}-{\omega_{15}}+{\omega_{16}}}\right),$
${{\tilde{\omega}}_{15}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}-{\omega_{3}}+{\omega_{4}}+{\omega_{5}}-{\omega_{6}}-{\omega_{7}}+{\omega_{8}}+{\omega_{9}}-{\omega_{10}}-{\omega_{11}}+{\omega_{12}}+{\omega_{13}}-{\omega_{14}}+{\omega_{15}}-{\omega_{16}}}\right),$
${{\tilde{\omega}}_{16}}=\frac{1}{4}\left({-{\omega_{1}}+{\omega_{2}}-{\omega_{3}}+{\omega_{4}}+{\omega_{5}}-{\omega_{6}}-{\omega_{7}}+{\omega_{8}}-{\omega_{9}}+{\omega_{10}}+{\omega_{11}}-{\omega_{12}}-{\omega_{13}}+{\omega_{14}}-{\omega_{15}}+{\omega_{16}}}\right).$ (8)
@@
Similarly, the antisymmetric sixteen-vertex model is that satisfying
@@
It is straightforward to see from Eq. (8) that, the symmetric sixteen-vertex model is transformed into an even eight-vertex model (Eq. (3) of Ref. [69] shows this case), while the antisymmetric case is transformed into an odd subcase.
We will show in the next section that each Ising model considered can be mapped to a symmetric/antisymmetric sixteen-vertex model, and further transformed into an even/odd eight-vertex model by Eq. (8). We can find that each case satisfies either the even free-fermion condition [41, 42]
@@
or the odd free-fermion condition [43]

### v2

# Free-fermion models and the two-dimensional Ising models under the zero field and imaginary field $i(\pi/2){k_{B}}T$

@@ Abstract
Ising model is famous in condensed matter and statistical physics. In this work we present a free-fermion formulation of the two-dimensional classical Ising models on the honeycomb, triangular and Kagomé lattices. Each Ising model is studied in the cases of a zero field and of an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$. We employ the decorated lattice technique, star-triangle transformation and weak-graph expansion method to exactly map each Ising model in both cases into an eight-vertex model on the square lattice. The resulting vertex weights are shown to satisfy the free-fermion condition. In the zero field case, each Ising model is an even free-fermion model. In the case of the imaginary field, the Ising model on the honeycomb lattice is an even free-fermion model while the models on the triangular and Kagomé lattices are odd free-fermion models. We obtain the exact solution of the Kagomé lattice Ising model under the imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$, a result not previously reported in the literature. We also show that the frustrated Ising models on the triangular and Kagomé lattices in the imaginary field still exhibit a non-zero residual entropy.
@@ I Introduction
Ising model plays a crucial role in the research of condensed matter and statistical physics of lattice systems, especially in understanding the phase transition via statistical mechanics. It has a long history since 1920s Niss (2005). The one-dimensional Ising model was first solved by Ising himself Ising (1925). Kramers and Wannier introduced the transfer matrix method to rederive the exact solution for the one-dimensional case and to study the two-dimensional case Kramers and Wannier (1941a, b). Onsager was the first to present the exact solution of the two-dimensional Ising model on the square lattice without a magnetic field, using the transfer matrix method Onsager (1944). This famous result has since been obtained through various approaches Kaufman (1949); Nambu (1950); Kac and Ward (1952); Schultz et al. (1964); Hurst and Green (1960); Hurst (1966); Gibberd and Hurst (1967); Fisher (1966); Samuel (1980), and has made a great impact on exactly solvable lattice models Newell and Montroll (1953); Baxter (1982, 1995); Lieb (1997); Baxter (2010). Lee and Yang proposed the solution for the case in an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$ in their work of Lee-Yang zeros Lee and Yang (1952). This solution has also been rederived using a variety of different methods Baxter (1965, 1967); McCoy and Wu (1967); Marshall (1971); Merlini (1974); Gaaff (1974); Lin and Wu (1988). Since then, the mathematically exact results for Ising models have attracted considerable attention, demonstrating their elegance and significance. Two-dimensional Ising models on the honeycomb lattice Husimi and Syôzi (1950); Syôzi (1950); Houtappel (1950); Temperley (1950), the triangular lattice Wannier (1950, 1973); Newell (1950); Potts (1955); Nourhani et al. (2023), the Kagomé lattice Syôzi (1951); Kanô and Naya (1953), and the checkerboard lattice Giacomini (1985); Wu (1986) were also exactly solved. Although the Ising model in a non-zero field is generally a well-known unsolved problem, the case of an imaginary field $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$ introduced by Lee and Yang Lee and Yang (1952), is unique and holds significant interest.
@@
In this work we focus on the exact solutions of anisotropic Ising models on typical two-dimensional lattices, both in the zero field and in the imaginary field. The Hamiltonian of N spins $\left\{{{s_{i}}=\pm 1,{\rm{}}i=1,\cdots,N}\right\}$ reads
$H=\sum\limits_{\left\langle{ij}\right\rangle}{J_{ij}{s_{i}}{s_{j}}}-{H_{{\rm{ex}}}}\sum\limits_{i=1}^{N}{{s_{i}}}~,$ (1)
where the first sum $\sum_{\left\langle{ij}\right\rangle}{}$ is over all nearest-neighbours, $J_{ij}$ denotes the directional anisotropic interaction and the field $H_{{\rm{ex}}}$} can be 0 or $i({{\pi\mathord{\left/{\vphantom{\pi 2}}\right.\kern-1.2pt}2}}){k_{B}}T$. The approach we employ is the mapping into the free-fermion model Fan and Wu (1969, 1970); Wu and Kunz (2004); Baxter (1986); Melotti (2021), which was proposed in the vertex model problem.
@@
The vertex model is also an important and intriguing topic in statistical mechanics. Studies of vertex models such as the six-vertex Slater (1941); Rys (1963); Lieb (1967a, b); Wu (1967); Lieb (1967c, d); Sutherland (1967); Yang (1967); Sutherland et al. (1967); Wu (1968); Glasser (1969), eight-vertex Fan and Wu (1969, 1970); Sutherland (1970); Baxter (1971, 1972); Kadanoff and Wegner (1971); Baxter (1973a, b, c, 1978, 2004); Wu and Kunz (2004) and sixteen-vertex models Wu (1969a, 1970, 1972a, 1972b); Gaaff and Hijmans (1975); Assis (2017) are closely related to those of Ising models. The free-fermion models, which are solvable eight-vertex models that satisfy the free-fermion condition, can be exactly solved using various methods such as the S-matrix Fan and Wu (1969), the Pfaffian Fan and Wu (1970) and the transfer matrix methods Felderhof (1973a, b, c); Bazhanov and Stroganov (1985a, b, c). Typically, an eight-vertex model on the square lattice is defined as either the even or odd subcase of the general sixteen-vertex model Assis (2017), as shown in Fig. 1. In this paper we adopt the notation of vertex configurations in Ref. Wu (1972a) (see Fig. 1 there), and refer to the eight-vertex model of even (odd) subcase satisfying the free-fermion condition as even (odd) free-fermion model.
@@
There have been lots of studies connecting the Ising model with the vertex model Gaaff (1974); Lin and Wu (1988); Debauche et al. (1991); Kooiman and van Leeuwen (1992); Morita and Tanaka (1993); Krčmár et al. (2022). Ref. Gaaff (1974) introduced a rather simple and direct correspondence between the square lattice Ising spin states and the arrow configurations of the sixteen-vertex model. The value $\pm 1$ of an Ising spin is naturally mapped to the direction of the arrow settled on the corresponding site in the sixteen-vertex model. By further performing a transformation from the sixteen vertex weights to those in the even subcase, the author showed that the Ising model is equivalent with an even free-fermion model, both for the zero field case and the imginary field case. The present paper is inspired by the success of this work Gaaff (1974). The method of transforming the square lattice Ising model to the free-fermion model is extended and applied to the Ising models on the honeycomb, triangular and Kagomé lattices.
The structure of the paper is as follows. In Sec. II we briefly introduce the methods and techniques used for transforming the Ising models into the free-fermion models, specifically the decorated lattice technique Syôzi (1951); Fisher (1959); Syôzi (1972), the star-triangle transformation Syôzi (1951); Fisher (1959); Onsager (1971); Syôzi (1972); Au-Yang and Perk (1989) and the weak-graph expansion Nagle (1968); Wu (1969b, 1972a); Nagle (1974). In Sec. III we present the exact results of the Ising models on the honeycomb, triangular and Kagomé lattices. Each of these models is shown to be equivalent with a free-fermion model, both in cases under the zero field and the imaginary field. The exact solutions are then obtained in the free-fermion formulation. We summarize in Sec. IV.
II Methods
It is essential to introduce the methods and techniques for transformation of lattice systems and vertex weights, before we present the main results of the Ising models. The decorated lattice technique Syôzi (1951); Fisher (1959); Syôzi (1972) and the star-triangle transformation Baxter (1982); Syôzi (1951); Fisher (1959); Onsager (1971); Syôzi (1972); Au-Yang and Perk (1989) help us to obtain a lattice structure, which is convenient for translating the Boltzmann factors to the vertex weights of a sixteen-vertex model. The resulting sixteen-vertex model is then transformed into an even or odd eight-vertex model by the weak-graph expansion Nagle (1968); Wu (1969b, 1972a); Nagle (1974).
@@ II.1 Decorated lattice technique
Decorated lattice technique is suitable for the case when we would like to add a spin site between a pair of nearest-neighbours, keeping the partition function invariant. As shown in Fig. 2, the interaction $J$ is replaced by $\tilde{J}$. In our consideration, the additional spin $\tilde{s}$ is used to represent the direction of the arrow on the edge linking $s_{1}$ and $s_{2}$. This may help us to build up an effective vertex model formulation of the Ising problem. Consider the Boltzmann factors of this pair contributed to the partition function, which should be conserved after the decorated lattice transformation. That is
$e^{-\beta J{s_{1}}{s_{2}}}=A\sum\limits_{\tilde{s}=\pm 1}{{e^{-\beta{\tilde{J}}\tilde{s}\left({{s_{1}}+{s_{2}}}\right)}}}~.$ (2)
@@
$e^{-\beta J}=A\left({{e^{-2\beta{\tilde{J}}}}+{e^{2\beta{\tilde{J}}}}}\right),~e^{\beta J}=2A.$
Thus the interaction $\tilde{J}$ and the multiplier $A$ are determined by
$\cosh\left({2\beta{\tilde{J}}}\right)=e^{-2\beta J}~,$ (3)
@@
Figure 2: The diagrammatic representation of the decorated lattice technique. A spin $\tilde{s}$ is added and the interaction $J$ is replaced by $\tilde{J}$.
@@ II.2 Star-triangle transformation
Fig. 3 shows the star-triangle transformation. The triangle with internal interactions $J_{1}$, $J_{2}$ and $J_{3}$ is replaced by a “star” consisting of an additional spin site coupled to the three vertices with interactions $J^{\prime}_{1}$, $J^{\prime}_{2}$ and $J^{\prime}_{3}$. The partition function should be invariant after the transformation, i.e., the Boltzmann factors of this unit should be conserved
$e^{-\beta\left({{J_{1}}{s_{1}}{s_{2}}+{J_{2}}{s_{2}}{s_{3}}+{J_{3}}{s_{3}}{s_{1}}}\right)}=B\sum\limits_{\tilde{s}=\pm 1}{e^{-\beta\tilde{s}\left({J^{\prime}_{1}{s_{3}}+J^{\prime}_{2}{s_{1}}+J^{\prime}_{3}{s_{2}}}\right)}}~.$ (5)
@@
${e^{-\beta\left({{J_{1}}+{J_{2}}+{J_{3}}}\right)}}=B\left[{{e^{-\beta\left({J^{\prime}_{1}+J^{\prime}_{2}+J^{\prime}_{3}}\right)}}+{e^{-\beta\left({-J^{\prime}_{1}-J^{\prime}_{2}-J^{\prime}_{3}}\right)}}}\right],$
${e^{-\beta\left({{J_{1}}-{J_{2}}-{J_{3}}}\right)}}=B\left[{{e^{-\beta\left({-J^{\prime}_{1}+J^{\prime}_{2}+J^{\prime}_{3}}\right)}}+{e^{-\beta\left({J^{\prime}_{1}-J^{\prime}_{2}-J^{\prime}_{3}}\right)}}}\right],$
${e^{-\beta\left({-{J_{1}}+{J_{2}}-{J_{3}}}\right)}}=B\left[{{e^{-\beta\left({J^{\prime}_{1}-J^{\prime}_{2}+J^{\prime}_{3}}\right)}}+{e^{-\beta\left({-J^{\prime}_{1}+J^{\prime}_{2}-J^{\prime}_{3}}\right)}}}\right],$
${e^{-\beta\left({-{J_{1}}-{J_{2}}+{J_{3}}}\right)}}=B\left[{{e^{-\beta\left({J^{\prime}_{1}+J^{\prime}_{2}-J^{\prime}_{3}}\right)}}+{e^{-\beta\left({-J^{\prime}_{1}-J^{\prime}_{2}+J^{\prime}_{3}}\right)}}}\right].$
The interactions $J^{\prime}_{1}$, $J^{\prime}_{2}$ and $J^{\prime}_{3}$ and the multiplier $B$ can be determined from the equations above Baxter (1982)
$\cosh\left(2\beta{J^{\prime}_{1}}\right)=\cosh\left({2\beta{J_{2}}}\right)\cosh\left({2\beta{J_{3}}}\right)$
$~~~~~~~~~~~~~~~~~-\sinh\left({2\beta{J_{2}}}\right)\sinh\left({2\beta{J_{3}}}\right)\coth\left({2\beta{J_{1}}}\right),$
$\cosh\left(2\beta{J^{\prime}_{2}}\right)=\cosh\left({2\beta{J_{3}}}\right)\cosh\left({2\beta{J_{1}}}\right)$
$~~~~~~~~~~~~~~~~~-\sinh\left({2\beta{J_{3}}}\right)\sinh\left({2\beta{J_{1}}}\right)\coth\left({2\beta{J_{2}}}\right),$
$\cosh\left(2\beta{J^{\prime}_{3}}\right)=\cosh\left({2\beta{J_{1}}}\right)\cosh\left({2\beta{J_{2}}}\right)$
$~~~~~~~~~~~~~~~~~-\sinh\left({2\beta{J_{1}}}\right)\sinh\left({2\beta{J_{2}}}\right)\coth\left({2\beta{J_{3}}}\right),$ (6)
$B=$ $\frac{{\left({1-t_{1}^{2}}\right)\left({1-t_{2}^{2}}\right)\left({1-t_{3}^{2}}\right)}}{4}\times$
$\left[{\frac{{-\sinh\left({2\beta{J_{1}}}\right)\sinh\left({2\beta{J_{2}}}\right)\sinh\left({2\beta{J_{3}}}\right)}}{{2\left({1-{t_{1}}{t_{2}}{t_{3}}}\right)\left({{t_{2}}{t_{3}}-{t_{1}}}\right)\left({{t_{1}}{t_{3}}-{t_{2}}}\right)\left({{t_{1}}{t_{2}}-{t_{3}}}\right)}}}\right]^{{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2}},$ (7)
with ${t_{i}}=\tanh\left({\beta{J_{i}}}\right)(i=1,2,3)$. One can easily verify that the star-triangle transformation directly connects the partition function of the triangular lattice model with that of the honeycomb lattice model. This enables us to simplify the derivation of the results for the triangular lattice model.
Figure 3: The diagrammatic representation of the star-triangle transformation. A spin $\tilde{s}$ is added and the interactions $J_{1}$, $J_{2}$ and $J_{3}$ are replaced by $J^{\prime}_{1}$, $J^{\prime}_{2}$ and $J^{\prime}_{3}$.
@@ II.3 Weak-graph expansion
Weak-graph expansion is a useful tool in lattice statistical problems Nagle (1968, 1974). In this paper, we use the version of weak-graph expansion method proposed by Wu, which was originally designed for the even eight-vertex model Wu (1969b) and later extended to the sixteen-vertex model Wu (1972a). Applying the method to the sixteen-vertex model enables us to obtain a new set of vertex weights, with the partition function invariant. The details of the method are shown in Ref. Wu (1969b), and we only list the result for all sixteen rearranged weights, expressed via a transformation matrix $\mathbf{M}$ as $\left\{{{\tilde{\omega}}_{1}},{{\tilde{\omega}}_{2}},\ldots,{{\tilde{\omega}}_{16}}\right\}^{T}=\frac{1}{4}\mathbf{M}\left\{{{\omega}_{1}},{{\omega}_{2}},\ldots,{{\omega}_{16}}\right\}^{T}$, with
[…]
@@
Similarly, the antisymmetric sixteen-vertex model is one satisfying
@@
It is straightforward to see from Eq. (24) that, the symmetric sixteen-vertex model is transformed into an even eight-vertex model (Eq. (3) of Ref. Wu (1972a) shows this case), while the antisymmetric case is transformed into an odd subcase.
We will show in the next section that each Ising model considered can be mapped to a symmetric/antisymmetric sixteen-vertex model, and further transformed into an even/odd eight-vertex model by Eq. (24). We can find that each case satisfies either the even free-fermion condition Fan and Wu (1969, 1970)
@@
or the odd free-fermion condition Wu and Kunz (2004)